# SGDFuse: SAM-Guided Diffusion for High-Fidelity Infrared and Visible Image Fusion

### v1

@@ IV-B4 Results and Analysis on the RoadScene Dataset
IV-C Ablation Study
@@
TABLE III: Quantitative results of all fusion variants in the four ablation studies. Bold numbers indicate the best performance.
@@ IV. Impact of HFAH
IV-C1 Impact of SAM
The Segment Anything Module (SAM) offers high-quality semantic masks in stage II, guiding the diffusion model for structural refinement and semantic enhancement, thereby improving image integrity and consistency. To assess SAM’s effectiveness, we conducted ablation studies by removing or modifying its components, designing three variants to analyze the impact of semantic guidance on fusion performance.
@@ IV-C1 Impact of SAM
Contrary to the previous setting, this variant retains the infrared semantic mask while removing the visible light mask, thereby evaluating the importance of the rich texture and edge information from the visible images in detail preservation and target recognition.
@@
IV-C2 Analysis of the Impact of the Number of MSFEM and TB Modules
To explore the optimal stacking depth for feature extraction, we performed ablation experiments by varying the number of MSFEM and TB modules in Stage I from 2 to 4 to investigate the influence of stacking depth on
@@ IV-C2 Analysis of the Impact of the Number of MSFEM and TB Modules
fusion performance. As shown in Fig. 10, increasing the number of modules generally improves fusion quality by enhancing detail restoration and structural consistency. However, performance saturates or slightly declines beyond three layers due to potential redundancy and overfitting. Therefore, we adopt three MSFEM and three TB layers as the final configuration, striking a balance between accuracy and efficiency, and providing a strong feature foundation for stage II diffusion optimization.
IV-C3 Impact of Two-Stage Training
IV-C4 Impact of the Diffusion Process
IV-C5 Impact of HFAH
IV-D Extension to high-level vision task
IV-D1 Object Detection
Fig. 11 shows that fused images significantly improve detection accuracy over single-modality inputs. Our method outperforms mainstream fusion approaches in both “Person” and “Car” categories, demonstrating superior object perception and generalization.
TABLE IV: Quantitative comparison of different fusion methods across categories. Best and second-best scores are marked in red and blue, respectively.
Category U2Fusion TarDAL PIAFusion IGNet CDDFuse Dif-fusion PFCFuse SHIP EMMA Text-Difuse DAFusion MaeFuse SAGE Ours
Background 0.976 0.969 0.972 0.971 0.975 0.976 0.975 0.976 0.977 0.972 0.977 0.975 0.981 0.979
Car 0.857 0.867 0.871 0.869 0.872 0.871 0.873 0.871 0.862 0.873 0.868 0.871 0.879 0.879
Person 0.682 0.681 0.680 0.679 0.685 0.684 0.683 0.681 0.656 0.673 0.679 0.685 0.675 0.686
Curve 0.703 0.755 0.721 0.769 0.789 0.774 0.737 0.745 0.735 0.743 0.756 0.774 0.747 0.780
Color Cone 0.725 0.783 0.766 0.779 0.794 0.772 0.748 0.760 0.769 0.764 0.778 0.796 0.783 0.798
Bike 0.716 0.769 0.739 0.754 0.785 0.786 0.729 0.742 0.638 0.748 0.750 0.787 0.774 0.789
mIoU 0.776 0.804 0.791 0.804 0.817 0.810 0.791 0.796 0.773 0.796 0.802 0.814 0.806 0.819
@@ IV-D2 Semantic Segmentation
V CONCLUSION

### v2

@@ 4.2.4 Results and Analysis on the RoadScene Dataset
4.3 Efficiency Comparison
To comprehensively compare with other state-of-the-art fusion methods, we evaluated the runtime efficiency of each approach in terms of computational cost and parameter scale. The detailed results are presented in Fig. 9. Computational cost is measured by the number of floating-point operations (FLOPs) required to process a single image, using the MSRS dataset for testing. TarDAL achieves higher efficiency due to its lightweight architecture with fewer parameters and lower computational cost; however, its fusion performance in terms of visual quality is inferior to our method. Diffusion-based methods introduce diffusion models, which significantly increase both parameter count and computational burden. SAGE incorporates SAM-generated segmentation masks as prompts to enhance semantic awareness in fusion but at the expense of increased computational complexity. Text-DiFuse employs a text-modulated diffusion model, improving flexibility and representational capacity in fusion, yet this also leads to higher computational overhead. In contrast, our method achieves a favorable balance by maintaining high-quality fusion results with moderate parameter size and computational cost, thus delivering better runtime efficiency. Overall, the analysis demonstrates that our approach strikes an effective trade-off between fusion performance and computational resource consumption, highlighting strong competitiveness and practical applicability.
4.4 Ablation Study
@@
4.4.1 Impact of SAM
The Segment Anything Module (SAM) offers high-quality semantic masks in stage II, guiding the diffusion model for structural refinement and semantic enhancement, thereby improving image integrity and consistency. To assess SAM’s effectiveness, we conducted ablation studies by removing
Table 3: Quantitative results of all fusion variants in the four ablation studies. Bold numbers indicate the best performance.
@@ IV. Impact of HFAH
or modifying its components, designing three variants to analyze the impact of semantic guidance on fusion performance.
@@
Contrary to the previous setting, this
variant retains the infrared semantic mask while removing the visible light mask, thereby evaluating the importance of the rich texture and edge information from the visible images in detail preservation and target recognition.
@@
4.4.2 Analysis of the Impact of the Number of MSFEM and TB Modules
To explore the optimal stacking depth for feature extraction, we performed ablation experiments by varying the number of MSFEM and TB modules in Stage I from 2 to 4 to investigate the influence of stacking depth on fusion performance. As shown in Fig. 10, increasing the number of modules generally improves fusion quality by enhancing detail restoration and structural consistency. However, performance saturates or slightly declines beyond three layers due to potential redundancy and overfitting. Therefore, we adopt three MSFEM and three TB layers as the final configuration, striking a balance between accuracy and efficiency, and providing a strong feature foundation for stage II diffusion optimization.
4.4.3 Impact of Two-Stage Training
4.4.4 Impact of the Diffusion Process
4.4.5 Impact of HFAH
4.5 Extension to high-level vision task
4.5.1 Object Detection
Fig. 11 shows that fused images significantly improve detection accuracy over single-modality inputs. Our method outperforms mainstream
@@ 4.5.1 Object Detection
fusion approaches in both “Person” and “Car” categories, demonstrating superior object perception and generalization.
Table 4: Quantitative comparison of different fusion methods across categories. Best and second-best scores are marked in red and blue, respectively.
Category Background Car Person Curve Color Cone Bike mIoU
U2Fusion 0.976 0.857 0.682 0.703 0.725 0.716 0.776
TarDAL 0.969 0.867 0.681 0.755 0.783 0.769 0.804
PIAFusion 0.972 0.871 0.680 0.721 0.766 0.739 0.791
IGNet 0.971 0.869 0.679 0.769 0.779 0.754 0.804
CDDFuse 0.975 0.872 0.685 0.789 0.794 0.785 0.817
Dif-fusion 0.976 0.871 0.684 0.774 0.772 0.786 0.810
PFCFuse 0.975 0.873 0.683 0.737 0.748 0.729 0.791
SHIP 0.976 0.871 0.681 0.745 0.760 0.742 0.796
EMMA 0.977 0.862 0.656 0.735 0.769 0.638 0.773
Text-Difuse 0.972 0.873 0.673 0.743 0.764 0.748 0.796
DAFusion 0.977 0.868 0.679 0.756 0.778 0.750 0.802
MaeFuse 0.975 0.871 0.685 0.774 0.796 0.787 0.814
SAGE 0.981 0.879 0.675 0.747 0.783 0.774 0.806
Ours 0.979 0.879 0.686 0.780 0.798 0.789 0.819
@@ 4.5.2 Semantic Segmentation
5 DISCUSSION
This paper proposes SGDFuse, an innovative framework for infrared and visible image fusion, designed to address the prevalent semantic gap in current fusion methods. Our core contribution lies in establishing a novel fusion paradigm: leveraging prior knowledge from a large segmentation model (SAM) as explicit semantic guidance to constrain and optimize the generative process of a conditional diffusion model, thereby achieving high-fidelity synthesis of cross-modal information. Extensive quantitative experiments, along with evaluations on downstream tasks such as object detection and semantic segmentation, consistently demonstrate that the fused images generated by SGDFuse significantly surpass state-of-the-art methods in terms of information integrity, detail fidelity, and semantic consistency. This robustly validates the superiority and effectiveness of our proposed ”semantic-guided generation” framework.
We have carefully considered the trade-off between performance and efficiency in our framework. The introduction of SAM and a diffusion model in SGDFuse admittedly increases computational complexity. However, we argue that this design is a necessary approach to overcome the fundamental limitations of traditional methods. Because traditional methods cannot accurately distinguish between foreground targets and background textures, they often lead to the loss of critical thermal radiation information or detail distortion, which fundamentally compromises the quality of the fused information. In contrast, SGDFuse exchanges a manageable computational overhead for a substantial leap in the semantic fidelity of the fused information.
Despite the state-of-the-art performance achieved by SGDFuse, there remain avenues for further exploration that constitute our future research directions. First, since SAM’s pre-training data is primarily based on visible light images, its generalization capability to the infrared domain has room for improvement. In the future, the precision of the semantic guidance can be further enhanced through domain-adaptive fine-tuning or by incorporating more specialized infrared segmentation models. Second, the iterative nature of the diffusion model limits its application in real-time scenarios. Future work will focus on model lightweighting, for instance, through knowledge distillation or by designing efficient samplers, to develop a version suitable for real-time applications and thereby broaden its operational boundaries.
6 CONCLUSION
@@
Acknowledgments
This research was supported by the National Natural Science Foundation of China (62301105, 61772319, 62272281, 62002200, 62202268), Shandong Natural Science Foundation of China (ZR2020QF012 and ZR2021MF068), Yantai science and technology innovation development plan(2022JCYJ031).
Data Availability Statement
Data available on request from the authors.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.